Status Report

NASA ISS On-Orbit Status 24 August 2009

By SpaceRef Editor
August 25, 2009
Filed under , , ,
NASA ISS On-Orbit Status 24 August 2009

All ISS systems continue to function nominally, except those noted previously or below. Underway: Week 13 of Increment 20.

Upon wakeup (~2:00am EDT), FE-1 Mike Barratt, FE-4 Bob Thirsk & FE-5 Frank De Winne started a new recording round for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), logging data from their Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. Meanwhile, Tim Kopra continues his extended session. [To monitor the crewmember’s sleep/wake patterns and light exposure, Mike & Koichi wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

For Padalka, Romanenko & Barratt, it was another periodic session with the Russian biomedical routine assessments PZEh-MO-7/Calf Volume Measurement. [MO-7 Calf measurements (left leg only) are taken with the IZOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference pints, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures.]

The CDR undertook his fifth session of the Russian behavioral assessment MBI-20 TIPOLOGIA, setting up the workstation, connecting equipment, suiting up and launching the program on the RSK1 laptop. [The FE-3 assisted him in donning the electrode cap, preparing the head for the electrodes, applying electrode gel from the Neurolab-RM2 kit and taking photographs. Data were recorded on a PCMCIA memory card and downlinked via OCA comm. MBI-20 studies typological features of operator activity of the ISS crews in long-term space flight phases, with the subject using a cap with EEG (electroencephalogram) electrodes. The experiment, which records EEGs, consists of the Lscher test, “adaptive biological control” training, and the games Minesweeper and Tetris. The Lscher color diagnostic is a psychological test which measures a person’s psychophysical state, his/her ability to withstand stress, to perform and to communicate. It is believed to help uncover the cause of psychological stress, which can lead to physical symptoms. An EEG measures and records the electrical activity of the brain.]

Mike did another a LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System) Phase 2 sampling session, using the Gram+ cartridges, setting up the equipment, preparing media and taking samples by surface swabbing. [LOCAD uses small, thumb-sized “microfluidic” cartridges that are read by the experiment reader. The handheld device tests a new analysis technology by sampling for the presence of gram negative bacteria in the sample in about 15 minutes. Lab-on-a-Chip technology has an ever-expanding range of applications in the biotech industry. Chips are available (or in development) which can also detect yeast, mold, and gram positive bacteria, identify environmental contaminants, and perform quick health diagnostics in medical clinics. The technology has been used to swab the MERs (Mars Exploration Rovers) for planetary protection. With expanded testing on ISS, this compact technology has broad potential applications in space exploration–from monitoring environmental conditions to monitoring crew health. After today, there are nine sessions remaining to complete the planned science requirements.]

Barratt also serviced the CIR (Combustion Integrated Rack) to support continuation of CIR MDCA (Multi-user Drop Combustion Apparatus) test point activities. [Mike first opened the FCF (Fluids & Combustion Facility) upper door, replaced a manifold bottle (#2002, with 40% O2, 60% CO2) with another bottle (#2014, 40% O2, 60% N2) plus exchanged the adsorber cartridge with a fresh one, then closed the facility up again.]

FE-4 Thirsk broke out and set up the equipment for a clinical evaluation, the U.S. PHS (Periodic Health Status) w/o Blood Labs exam, with Frank De Winne assisting as CMO (Crew Medical Officer). For the examination, Bob used the AMP (Ambulatory Medical Pack) and automatic blood pressure cuff from the ALSP (Advanced Life Support Pack). Afterwards, the two crewmembers switched roles and De Winne became Subject, with Thirsk acting as CMO. [The second part of PHS, Subjective Clinical Evaluation, was also performed afterwards. All data were then logged on the MEC (Medical Equipment Computer) and the hardware stowed. The PHS exam, with PCBA analysis and clinical evaluation, is guided by special software (IFEP, In-Flight Examination Program) on the MEC laptop.]

CDR Padalka, with Frank De Winne assisting, spent an hour on the TVIS treadmill for the periodic Russian PZE-MO-3 test for physical fitness evaluation, using the TVIS in unmotorized (manual control) mode and wearing the Kardiokassette KK-2000 belt with three chest electrodes. [The fitness test, controlled from the RSE-Med laptop, yields ECG (electrocardiogram) readings to the KK-2000 data storage device, later downlinked via the Regul (BSR-TM) payload telemetry channel. Before the run, the KK-2000 was synchronized with the computer date/time readings. For the ECG, the crewmembers worked out on the treadmill, first walking 3 min. up to 3.5 km/h, then running at a slow pace of 5-6 km/h for 2 min, at moderate pace of 6.5 km/h, followed by the maximum pace not exceeding 10 km/h, then walking again at gradually decreasing pace.]

In the Kibo JPM (JEM Pressurized Module), working on the Saibo Rack, FE-2 Kopra reinstalled closeout cloth at a closeout panel (JPM1D7-08) as well as soft dummy panels at JPM1A5 & JPM1A2, taking photos afterwards of the final configuration.

In preparation for the transfer of the JRSRS2 Rack from the JPM to the HTV1 after its arrival, Kopra also retrieved, transferred and restowed the IHI Corp. K-BAR (Knee-Brace Assembly Replacement) units, left & right.

The FE-2 relocated a RAM (Radiation Area Monitor) unit from the JPM, loc. JPMA5_D2 (inside of door), to the U.S. Lab, loc. LAB1O5_D2 (center of inside of locker door).

Afterwards, Tim used a digital still camera to take documentary photographs of each individual locker of the LAB1O5 rack. [These photos were requested to give ground specialists insight into the configuration of the new rack location for CHeCS (Crew Health Care Systems) hardware and determine if the stowage plan for this rack is acceptable. This also assists with the stowage plan for CHeCS hardware arriving on flight 17A.]

Roman, with some assist from Gennady, had ~1.5 hrs set aside for the periodic RS (Russian Segment) window inspection & photography in the SM & DC1, using a tool kit with ruler, adhesive tape, 90-deg equilateral triangle & measuring tape, the NIKON D2 X digital camera with 28-70 mm lens, a flash attachment, and sketches of the windows under scrutiny (1, 3, 5, 12 in SM, plus the VP1 EVA hatch window in DC1) with previous detected flaws marked and flaw tables. [Purpose of the activity is to assess the condition of the window panes for deterioration as compared to the data from previous increments (appearance of new cavities, scratches, discolorations, or spots reducing transparency, or an increase in the size of old flaws), plus photography. Then images and data tables were stored on the RSK1 laptop for subsequent downlink via OCA.]

Using the Russian AK-1M sampler, the FE-3 collected the periodic air samples in SM & FGB.

Afterwards, Romanenko conducted another 30-min. session with the new ocean observations program, DZZ-13 “Seiner”, to obtain data on color field patterns and current cloud cover conditions over the Northern Atlantic between the coasts of Newfoundland and the Western Sahara. [The experiment uses visual observation, videography (HDV camcorder, PAL mode) and selective photography (NIKON D2X with AF-S Zoom-Nikkor 80-200 mm lens) of color-contrast images and large discontinuities in cloud fields along the flight path, controlled from the RSK-1 laptop. Roman’s photography had to be accompanied by a continuous non-stop video recording of underlying terrain using the HDV camera securely fixed above SM Window #8 precisely in nadir using the LIV adapter.]

Padalka had ~1 hr for V3 fan screen and air duct cleaning in the DC1 Docking Compartment.

The CDR also worked on Russian SUBA crew support laptops to update their HDDs (Hard Disk Drives) with new antivirus software. [After copying update scripts and files from the FS (File Server) onto a flash drive, Gennady loaded the HDDs of the RSK1, RSK2, RSE1, RSE2 and RSS2 laptops.]

In the Soyuz 18S spacecraft, Gennady performed another health check of the KhSA Cooler/Dehumidifier Assembly’s V1 fan in the DM (Descent Module) by turning the V1 fan on and the V2 fan off, then checking air flow. [On 6/25, a planned replacement of the apparently faulty fan in the Soyuz 18S DM with a new unit proved to be not necessary after Padalka configured a jumper bypass which successfully recovered functionality of the air conditioner fan. Today’s activity was to check up on the fix.]

Afterwards, the CDR had 2 hrs set aside to conduct the periodic lighting fixtures audit in the RS.

De Winne serviced the EHS TOCA (Environmental Health System/Total Organic Carbon Analyzer) by completing the standard changeout of its WWB (Waste Water Bag).

Frank also performed the periodic WPA (Water Processor Assembly) sample analysis in the TOCA, after first priming (filling) the TOCA water sample hose. After the approximately 2 hr TOCA analysis, results were transferred to SSC-7 (Station Support Computer 7) via USB drive for downlink, and the data were also logged for calldown. [The current procedure is a work-around for TOCA’s failed catalyst.]

In the Lab, the FE-5 prepacked the supply tank of the FSS (Fluid System Servicer), preparing it for return on 17A.

Afterwards, Frank had ~2 hrs for clearing the aisle ways in the USOS (US Segment), including JPM, for rack translations during the 17A mission, stowing any remaining equipment or cargo bags that would impede a 50”x50” path as will be required for rack translation.

After setting up the camcorder for recording video, FE-4 Thirsk, with the FE-5 lending a hand, reviewed instructions and then conducted the EPO (Educational Program Operation) discussion & demo of “Conservation of Linear Momentum Demo”. Afterwards, Bob recorded a second EPO demo, “Time”, discussing time zones and how time is defined both on Earth and in space. The hardware was then restowed. [The demos were timed such that EPO ground personnel could provide real-time feedback during the demo.]

Working in the ESA COL (Columbus Orbital Laboratory), Bob Thirsk re-arranged EPM (European Physiology Module) items and consolidated the storage configuration for the EPM and ETC (European Transport Carrier) stowage locations.

The FE-4 set up and reactivated the ALTEA (Anomalous Long Term Effects on Astronauts) dosimeter unit which had been turned off and moved aside for access to the OGA (Oxygen Generator Assembly) rack.

For the MPLM (Multi-Purpose Logistics Module) arrival, Tim Kopra reviewed OBT (On Board Training) material on CBM (Common Berthing Mechanism) procedures.

The FE-2 removed and prepacked 11 nonfunctional SPS LHAs (Secondary Power System Lamp Housing Assemblies) for return on 17A, in preparation for spares to be scavenged from the MPLM “Leonardo” and to be brought up on the subsequent HTV-1 mission.

Tim also had an hour set aside for regular crew departure preparations, working on the standard end-of-increment cleanup preparatory to his return to Earth on STS-128/17A. [It is usual for crewmembers to be granted reduced workdays for making their departure preparations, as their return date approaches.]

FE-3 Romanenko completed another periodic inventory/audit of the Russian medical cabinet supplies.

In the U.S. A/L (Airlock), Mike Barratt performed the yearly maintenance on the EMU (Extravehicular Mobility Unit) #3006 water tanks. In addition, to prepare for a potential EVA in support of HTV docking, the FE-1 filled his LCVG (Liquid Cooling & Ventilation Garment).

Barratt also reviewed a video with a refresher course for working with the HMS CMRS (Health Maintenance System/Crew Medical Restraint System).

The FE-3 did the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Roman also completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The crew performed their regular daily 2.5-hr. physical workout program on the CEVIS cycle ergometer (FE-4), TVIS treadmill with vibration isolation (CDR, FE-1, FE-2, FE-3, FE-5), ARED advanced resistive exercise device (FE-1, FE-2, FE-3, FE-4, FE-5), and VELO cycle ergometer with bungee cord load trainer (CDR).

Afterwards, Mike transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

The FE-1, FE-2, FE-4 & FE-5 had their periodic PMCs (Private Medical Conferences), via S- & Ku-band audio/video, Frank at ~7:45am, Bob at ~9:40am, Tim at ~10:55am, Mike at ~2:05pm EDT.

Barratt & Kopra also had their weekly PFCs (Private Family Conferences), via S-band/audio and u-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Mike at ~8:20am, Timothy at ~4:15pm.

OGS Status Update: On 8/21 (Friday) the crew successfully replaced the suspect Water ORU (Orbital Replacement Unit) with a new unit. The OGS (Oxygen Generator System) was started successfully on the same day, and the delta pressure across the pump ORU came up with a nominal value and has remained so over the weekend (currently running at ~7.8 psid). The OGS is now considered to be operating nominally.

No CEO (Crew Earth Observation) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

Significant Events Ahead (all dates Eastern Time, some changes possible!):
08/25/09 — STS-128/Discovery/17A launch – MPLM (P), LMC (~1:36am EDT)
09/06/09 — STS-128/Discovery/17A landing (KSC; ~8:38pm)
09/10/09 — H-IIB (JAXA HTV-1) launch (~1:04pm EDT)
09/16/09 — H-IIB (JAXA HTV-1) berth w/SSRMS
09/29/09 — Progress 34P undock
09/30/09 — Soyuz TMA-16/20S launch
10/02/09 — Soyuz TMA-16/20S docking (SM aft, until MRM-2 w/new port)
10/11/09 — Soyuz TMA-14/18S undock
10/14/09 — H-IIB (JAXA HTV-1) unberth
10/15/09 — Progress 35P launch
11/10/09 — 5R/MRM-2 (Russian Mini Research Module 2) on Soyuz-U
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/07/09 — Soyuz TMA-17/21S launch
12/26/09 — Progress 36P launch
02/03/10 — Progress 37P launch
02/04/10 — STS-130/Endeavour/20A – Node-3 + Cupola
03/18/10 — STS-131/Discovery/19A – MPLM(P), LMC
04/02/10 — Soyuz TMA-18/22S launch
04/27/10 — Progress 38P launch
05/14/10 — STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1
05/29/10 — Soyuz TMA-19/23S launch
06/25/10 — Progress 39P launch
07/29/10 — STS-133/Endeavour (ULF5 – ELC4, MPLM) or STS-134/Discovery (ULF6 – ELC3, AMS)
08/11/10 — Progress 40P launch
09/16/10 — STS-133/Endeavour (ULF5 – ELC4, MPLM) or STS-134/Discovery (ULF6 – ELC3, AMS)
09/29/10 — Soyuz TMA-20/24S launch
10/19/10 — Progress 41P launch
11/??/10 — ATV2 – Ariane 5 (ESA)
12/??/11 — 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton

SpaceRef staff editor.